OpenStudy (waheguru):

what is teh equation if m=4/3 and the coordinates are (-2,0) HELPPPPPPPPPPPPPP

OpenStudy (amistre64):

youre given a point and a slope; use point slope form

OpenStudy (cathyangs):

ok. so m is the variable usually used for slope, and if that's the case, the equation is y=4/3x+-8/3, which you can find by plugging the x and y coordinates given.

OpenStudy (waheguru):

i got ann answer but its not matching with the answers at the back

OpenStudy (cathyangs):

or you can use point slope.

OpenStudy (amistre64):

by saying "what is the equation" is a bit ambiguous since there are many forms of the same equation

OpenStudy (waheguru):

g got y=4/3x-8/3?????

OpenStudy (waheguru):

but insted of minsus its plus in the answers?

OpenStudy (anonymous):

Using slope point form: y-y1=m(x-x1) y-0=4/3(x+2) Y=4x/3+2 2Y=4x+6 4x-2y+6=0

OpenStudy (waheguru):

what?

OpenStudy (yash2651995):

ans at back can be wrong too.. remember "Don't blindly trust a book, a misprint in medical books can be lethal" -my words :P

OpenStudy (waheguru):

just find the equation for the line y=mx+b

hero (hero):

Aroub, help him

OpenStudy (aroub):

yea i can help! but as i can see.. everyone IS helping him!! :)

hero (hero):

Aroub, get your hands dirty and help out like everyone else :P

OpenStudy (cathyangs):

SORRY!!!!! It's y=4/3x+8/3...I forgot the negative. DX